OpenStudy (haleyelizabeth2017):

Honors Algebra 2, continued.

OpenStudy (haleyelizabeth2017):

divide and simplify if possible \(\large\frac{ \sqrt {250x^{16}} }{ \sqrt {2x} }\)

OpenStudy (ericwise):